# Sequential Diagnosis with Language Models

### v1

@@ Abstract
11footnotetext: Equal contribution
@@ 3.2 MAI Diagnostic Orchestrator
After internal deliberation, the panel reaches consensus on one of three actions: asking questions, ordering tests, or committing to a diagnosis (if certainty exceeds threshold). Before tests are ordered, an optional budget tracker can be invoked to estimate both the cumulative medical costs so far and the cost of each test in the order.
@@
  • No Budget – Full panel with no explicit cost tracking or budget limitations.
@@ 3.3 Physicians
To establish human performance, we recruited $21$ physicians practicing in the US or UK to act as diagnostic agents. Participants had a median of 12 years [IQR 6-24 years] of experience: 17 were primary care physicians and four were in-hospital generalists. Each physician received the same initial vignette as the AI agents, and interacted with an identical Gatekeeper interface. No limits were imposed on session duration or the number of tests ordered. Cases were drawn from the hidden test set and case order was randomized for each participant to mitigate ordering effects. Physicians were explicitly instructed not to use external resources, including search engines (e.g., Google, Bing), language models (e.g., ChatGPT, Gemini, Copilot, etc), or other online sources of medical information. They were blinded to the correctness of their diagnosis, and were asked to complete as many as possible during the study period.
@@ 4 Results
Figure 7: Pareto-frontier showing diagnostic accuracy versus average cumulative monetary cost for each agent. Off-the-shelf models were evaluated using a uniform baseline prompt (see Figure 4). MAI-DxO, built on top of the o3 model, achieves Pareto dominance over both off-the-shelf models and practicing physicians.
@@ Acknowledgments
We are grateful to NEJM Group for permission to use the NEJM cases in the research reported in this paper. The research described here has benefited from work across many teams at Microsoft. We thank esteemed colleagues both inside and outside of Microsoft for sharing their insights including Bryan Bunning, Nando de Freitas, Kenneth D Mandl, Andrija Milicevic, Joseph Petro, Hoifung Poon, David Rhew, Karén Simonyan, Eric Topol, and Jim Weinstein. Gianluca Fontana and Kevin Hawkins (Prova Health) provided support on the health economics and outcomes section.

### v2

@@ Abstract
11footnotetext: Equal contribution00footnotetext: Correspondence: hanori@microsoft.com, horvitz@microsoft.com
@@ 3.2 MAI Diagnostic Orchestrator
After a “Chain of Debate” – a structured deliberation between sets of agents with defined roles – the panel reaches consensus on one of three actions: asking questions, ordering tests, or committing to a diagnosis (if certainty exceeds threshold). Before tests are ordered, an optional budget tracker can be invoked to estimate both the cumulative medical costs so far and the cost of each test in the order.
@@
  • Unconstrained Budget – Full panel with no explicit cost tracking or budget limitations.
@@ 3.3 Physicians
To establish human performance, we recruited $21$ physicians practicing in the US or UK to act as diagnostic agents. Participants had a median of 12 years [IQR 6-24 years] of experience: 17 were primary care physicians and four were in-hospital generalists. Each physician received the same initial vignette as the AI agents, and interacted with an identical Gatekeeper interface. No limits were imposed on session duration or the number of tests ordered. Cases were drawn from the hidden test set and case order was randomized for each participant to mitigate ordering effects. Physicians were blinded to the correctness of their diagnosis, and were asked to complete as many as possible during the study period.
Physicians were explicitly instructed not to use external resources, including search engines (e.g., Google, Bing), language models (e.g., ChatGPT, Gemini, Copilot, etc), or other online sources of medical information. Although limiting the use of search engines may not accurately reflect physicians’ real world clinical practice, the original NEJM cases are accessible online, and we sought to prevent participants from readily obtaining correct answers through external searches. Additionally, certain search engines offer AI-generated summaries, potentially providing diagnostic hints. By restricting physicians’ access to language models, we aimed specifically to assess their intrinsic diagnostic capabilities, rather than indirectly evaluating the performance of available generative artificial intelligence tools.
@@ 4 Results
Figure 7: Pareto-frontier showing diagnostic accuracy versus average cumulative medical cost for each agent. Each test ordered on the diagnostic journey adds to total expenditure. Off-the-shelf models were evaluated using a uniform baseline prompt (see Figure 4). MAI-DxO, built on top of the o3 model, achieves Pareto dominance over both off-the-shelf models and practicing physicians – that is, at any price point, MAI-DxO achieves higher accuracy over other solutions.
@@ Acknowledgments
We are grateful to NEJM Group for permission to use the NEJM cases in the research reported in this paper. The research described here has benefited from work across many teams at Microsoft. We thank esteemed colleagues both inside and outside of Microsoft for sharing their insights including Bryan Bunning, Nando de Freitas, Kenneth D Mandl, Andrija Milicevic, Joseph Petro, Hoifung Poon, David Rhew, Adam Rodman, Nigam Shah, Ted Shortliffe, Karén Simonyan, Eric Topol, Bob Wachter, and Jim Weinstein. Gianluca Fontana and Kevin Hawkins (Prova Health) provided support on the health economics and outcomes section.
@@ Appendix B Visualizing Model and Clinician Performance
Appendix C Recent CPC Case
The most recently published CPC case, “Case 18-2025: A 63-Year-Old Woman with Dyspnea on Exertion”, was published on June 26, 2025.
MAI-DxO was run on the case on June 26, 2025 and the transcript between MAI-DxO and the Gatekeeper for the case is shown on the pages below. Note that the transcript shows indicators of which test results were produced from the case text, and which ones were generated synthetically for illustrative reasons – this information is not revealed in the interface used by models and physicians attempting to solve the case.